TETSXPREME:

7^8 divide about 7^4

Tranquility:

\(\Large \dfrac {a^b}{a^c} = a^{b-c}\)

TETSXPREME:

so i do 8-4?

Tranquility:

\(\dfrac{7^8}{7^4} = ??\)

Tranquility:

@tetsxpreme wrote:
so i do 8-4?
Yes, to find the exponent

TETSXPREME:

7^4?

TETSXPREME:

i think i get it now

Tranquility:

7^4 is correct!

TETSXPREME:

thank you